write the equation of a line that passes through the point(-1,3)and is parallel to the graph of y=2x-6

Answer by robertb(5830)

The slope of the line has to be 2. By the point-slope form, the equation is y-3 = 2(x--1).
y-3 = 2(x+1),
y = 2x+2+3,
y = 2x+5.